OpenStudy (anonymous):

Find the inverse of f(x)=3e^(x+2/4)-5

OpenStudy (anonymous):

first off \[\frac{2}{4}=\frac{1}{2}\] is there a typo here?

OpenStudy (anonymous):

its x+2 over 4

OpenStudy (anonymous):

oh wait, i bet it is \[f(x)=e^{\frac{x+2}{4}}-5\]

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

but a 3 in front of e...

OpenStudy (anonymous):

ok we write \[y=3e^{\frac{x+2}{4}}-5\] and solve for x

OpenStudy (anonymous):

\[\frac{y}{3}=e^{\frac{x+2}{4}}-5\] \[\frac{y}{3}+5=e^{\frac{x+2}{4}}\]

OpenStudy (anonymous):

do i need to use: e^A=B

OpenStudy (anonymous):

\[\frac{x+2}{4}=\ln(\frac{y+15}{3})\]

OpenStudy (anonymous):

A=Ln(B)

OpenStudy (anonymous):

\[x+2=4\ln(\frac{y+15}{3})\]

OpenStudy (anonymous):

\[x=4\ln(\frac{y+15}{3})-2\]

OpenStudy (anonymous):

\[f^{-1}(x)=4\ln(\frac{x+15}{3})-2\]

OpenStudy (anonymous):

how did you get y+15?

OpenStudy (anonymous):

\[\frac{y}{3}+5=\frac{y+15}{3}\] by adding the fractions

OpenStudy (anonymous):

d'oh! gotcha. Thanks a million